At 3.36 percent interest, how long does it take (in years) to quadruple your money? Answer to two decimals.
Answer #1

Let X be the amount invested ie; present value

Future value = Present value * (1+ rate per period)^ no. of periods

4X =X*1.0336^n

1.0336^n = 4

n = log 4 / log 1.0336

= 0.60205999132/0.01435250065

= 41.9480901623

= 41.95 years